You want to be a millionaire when you retire in 40 years. You can much more will you have to save each month if you wait 10 y Please show formulas, calculations, and steps! Show all work
Answer #1

Number of periods = 40*12 = 480
Rate per month = 12.5%/12
FV = 1,000,000
Monthly Payment = FV/(1-(1+12.5%/12)-480)/(12.5%/12) =  1,000,000/(1-(1+12.5%/12)-480)/(12.5%/12) = 72.53

New Number of Periods = 30*12 = 360
Monthly Payment = FV/(1-(1+12.5%/12)-360)/(12.5%/12) =  1,000,000/(1-(1+12.5%/12)-360​​​​​​​)/(12.5%/12) = 255.91

More savings per month required = 255.91 - 72.53 = 183.38 (Option e is correct option)

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well